OpenStudy (anonymous):

variable z varies jointly with x and y. x=-12, y=1/3 , z=30. whats the equation

OpenStudy (radar):

2x-xy z=--------- 2y

OpenStudy (amistre64):

z = kxy 30 = k(-12)(1/3)

OpenStudy (amistre64):

30*3 ---- = k -12

OpenStudy (amistre64):

k = -5 right?

OpenStudy (amistre64):

z = -5xy

OpenStudy (amistre64):

30*3 not equal 60 .... forgot how to multiply

OpenStudy (amistre64):

z = -7.5 xy